OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 2x - one over four (12x + 4) = 3 Step 1: 2x + 3x + 1 = 3 Step 2: 5x + 1 = 3 Step 3: 5x + 1 - 1 = 3 - 1 Step 4: 5x = 2 In which step was the error made and why? Step 1, because incorrect signs have been used Step 2, because the terms have been added incorrectly Step 3, because a number has been subtracted from both sides of the equation No error was made

OpenStudy (armyrangers):

Ok witch one do you think it is

OpenStudy (anonymous):

a or c

OpenStudy (anonymous):

what do you think @ARMYRANGERS

OpenStudy (armyrangers):

Well am not aloud to just tell you the answer but I will tell you one of those that you have are right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a

OpenStudy (aum):

\(\Large 2x - \frac 14(12x+4) = 3 \\ \Large 2x - 3x - 1 = 3 \)

OpenStudy (armyrangers):

Ok well there is your answer

OpenStudy (anonymous):

ty